# Minimizing the number of edges in LC-equivalent graph states

### v1

@@ IV Creation of generalized repeater graph states
To demonstrate the utility of our algorithms, we consider the generation of photonic repeater graph states for long distance quantum communication. Repeater graph states (RGS) [6] are a key component of all-photonic quantum repeaters, enabling the distribution of a single unit of bipartite entanglement, known as an ebit over long distances. The implementation of such all-photonic repeaters has been extensively studied [26, 52, 53, 54] but often comes with a large overhead in the number of required single photon sources. A more general class of all-photonic states called generalized repeater graph states (gRGS) was introduced in Ref. [11], which offers a better scaling between photon overhead and the number of shared ebits. Here, we apply our MER algorithms to further minimize the number of single photons and optical fusion operations for the creation of gRGS.
@@
Figure 4: (a) and (b) show the required number of resource states and fusions, respectively, for creating a single instance of $G$+$L$ gRGS with 64 edges and 36 vertices (see Fig. 3(b)). We compare the naive approach of direct gRGS construction (blue curve) with the Commute-LC protocol (orange curve). The Commute-LC protocol requires nearly two orders of magnitude fewer resources than the naive approach, especially at lower fusion success probabilities. The resource savings decrease to less than an order of magnitude for relatively higher success probabilities.
@@
The probabilistic nature of fusions necessitates a significant amount of resources to construct practically useful graph states. Creating all-photonic graph states involves multiple rounds of fusions starting with resource states that are fused to create intermediate graph states, which are incrementally fused to construct the final graph state. For a given graph, Ref. [28] constructs a so-called fusion network, indicating the order in which fusions should occur. In general, it is desirable to fuse two graph states of similar sizes to mitigate the effects caused by fusion failure. Moreover, multiple fusion networks may exist for a given graph, each potentially requiring different amounts of resources. Ref. [28] addresses these challenges and reduces the resource overhead for photonic graph state creation by optimizing (1) the number of required fusions, (2) the number of resource states, and (3) the fusion order. Moreover, they provide a numerical tool for this purpose called the OptGraphState.
@@
Here, we exploit the structure of generalized repeater graph states to optimize their experimental creation. Such graphs consist of a densely-connected central graph ($G$) and leaves ($L$) that are attached to each vertex of $G$. We find an approximate MER ($H$) of the central graph $G$ using our SA algorithm. Naively, one could fuse resource states to create $G$+$L$ directly. However, a more resource-efficient method for creating a gRGS could be: (1) creating the graph $H$ using fusions, (2) applying LC-gates to convert the graph $H$ to $G$, and (3) adding leaves $L$ to $G$. We will refer to this way of construction as Intermediate-LC. However, the addition of leaves at the end constrains the order of fusions which could affect the probability of creation of the gRGS. We propose commuting the LC-gates to the beginning of the protocol that lifts this constraint and allows for an optimal fusion order. We refer to this protocol as Commute-LC. In conclusion, the Commute-LC protocol has two steps: (1) apply the LC-gates to the resource states at the beginning of the protocol, (2) follow fusion order to create the graph $H$+$L$ using fusions. The LC-gates here correspond to the conversion of the graph $H$ to $G$. We should note that fusions involve the application of single-qubit Pauli correction operations, but they can be commuted through Clifford gates since these are Pauli corrections.
@@
We illustrate the construction of $G$+$L$ using the Commute-LC protocol in Fig. 3(a). Following the steps in Commute-LC, we first apply LC-gates to the appropriate qubits, followed by fusing states according to the order obtained for creating $H$+$L$. Applying LC-gates converts the resource states into LC-equivalent non-graph states. The Commute-LC protocol fuses these non-graph states following the order of $H$+$L$ to obtain the desired graph state $G$+$L$. Appendix D details how to determine locations of LC-gates and how following the order for $H$+$L$ constructs the $G$+$L$ graph. We note here that Ref. [28] also applies LC-gates to their qubits during their optimization. However, these LC-gates only convert a fully-connected subgraph (clique) into a star. In contrast, by applying SA to the central graph $G$ in Protocols 1 and 2, we generalize the use of LC-gates to simplify the fusion process not just for cliques but for arbitrary densely connected subgraphs.
@@
In Fig. 4, we compare two methods: (1) the naive approach of creating $G$+$L$ directly, and (2) creation following the Commute-LC protocol. We use the Python package OptGraphState to estimate the required resources. We consider a gRGS that can share 3 ebits across a network  [^2] Similar to Ref. [11], shown in Fig. 3(b). This gRGS has a central graph ${G}$ with 18 vertices and 46 edges, shown in Fig. 3(b). We find an approximate MER of this graph with 23 edges using the SA algorithm (with $k_{\textrm{max}}=10000$ and $T(1)=100$). We observe that commuting LC-gates to the beginning and following fusion order for creating $H$+$L$ significantly reduces the required resources. This reduction is nearly two orders of magnitude for fusions in high-loss conditions.
V Outlook
@@
We have proposed three algorithms to address the edge-minimization problem. First, a simulated annealing (SA) approach efficiently finds approximate MERs for graphs with up to 100 qubits. Second, we introduce an integer linear programming (ILP) formulation that yields exact solutions, albeit with potentially exponential runtime. Third, we combine both methods into a hybrid SA+ILP algorithm, enabling exact MER identification for graphs with up to 16 qubits.
@@
We benchmark all algorithms on two graph models: bounded-degree and Erdős-Rényi. Additionally, we extend the ILP to handle weighted-edge minimization and prove that this variant is NP-complete. Importantly, we show that the scalable SA approach manages to find approximate MERs with less than 10% more edges on average than the actual MERs for a large set of randomly sampled graphs.
@@
We have applied our SA algorithm to minimize the required resources for all-photonic generalized repeater graph states (gRGS) demonstrating its practical utility for long distance quantum communication. Specifically, we have proposed to apply LC-gates corresponding to local complementation to the resource states before fusing them to create the gRGS states. We then demonstrate that SA can be combined with existing methods from Ref. [28] to reduce the required number of single-photon sources and fusion operations be more than an order of magnitude for realistic lossy implementations.
@@
Our results opens up a number of future directions. Finding the MER can reduce the number of required CZ gates for creating a desired graph state between qubits following the idea in Ref. [31]. However, there are examples of graph states where the minimum number of edges up to local complementations is greater than the minimum number of two-qubit gates needed to create the state [57, 58, 59]. It would be interesting to further investigate the exact connection between MERs and the minimal number of CZ-gates for generating qubit graph state generation
@@
Furthermore, we have focused exclusively on transformations between graph states that are allowed with local Clifford unitaries. In practical scenarios, however, one might consider arbitrary local unitaries (LU), which define a strictly coarser equivalence relation on graphs than the LC-equivalence relation, as was first shown in [60]. Graph states under local unitary equivalence are significantly less well-understood, but this topic has seen recent investigations [61, 62, 63], closely relating this equivalence to the Clifford hierarchy. One natural follow-up question is whether edge minimization of graphs under local unitaries can similarly be phrased as an ILP (using, for example, the algorithm from [63]), and, if so, to what extent this coarser equivalence can reduce the edge count.
Acknowledgements.
@@
We thank Tim Coopmans and Sebastiaan Brand for useful conversations, and Luise Prielinger for her help with the HPC. JH acknowledges funding from the Dutch Research Council (NWO) through a Veni grant (grant No.VI.Veni.222.331). JH and HS acknowledge funding from the Quantum Software Consortium (NWO Zwaartekracht Grant No.024.003.037). JB acknowledges support from The AWS Quantum Discovery Fund at the Harvard Quantum Initiative. Part of this work was performed while HS was on a research visit at Harvard University funded by a Quantum Delta NL travel grant.
@@
A release of our code is publicly available at https://doi.org/10.5281/zenodo.15534419 and the accompanying data at https://doi.org/10.5281/zenodo.15534839, with the expectation that it will serve as a valuable resource for the quantum information community in addressing practical optimization challenges related to graph states. The usage guidelines and notes can be found at the GitHub repository https://github.com/arr0w-hs/graph_state_optimization.
@@ Appendix A LC-equivalence algorithm for graph states
$M^{T}\Omega M=\Omega~{}\textrm{, ~{}where }\Omega=\begin{bmatrix}0&I\\ I&0\\ \end{bmatrix}\ ,$ (6)
@@
The generator sets of the two graph states |$\ket{G}$ and |$\ket{H}$ under consideration are given by
@@
Eq. (12) is a set of $n^{2}$ equations in $4n$ variables $(p_{i},q_{i},r_{i},s_{i})$ with a constraint (Eq. (10)) that is non-linear, so in principle not easily solvable. However, through an ingenious argument, Bouchet [36] proved that only a small number of solutions to Eq. 12 have to be checked to verify if there exists one that also satisfies Eq. (10), making the total runtime polynomial [36]. However for edge minimization, we will ignore this extra structure and just use Eqs. 12 and 10 to construct our ILPs.
Appendix B Derivation of ILP
In this section, we show how we formulate an ILP for edge minimization inspired by Bouchet’s algorithm.
Let us first formally state our optimization problem. The input is a single graph $G$ and the output is an LC-equivalent graph $H$ with a minimal number of edges. Mathematically it is framed as follows:
@@ Appendix B Derivation of ILP
We now utilize the formulation of LC-equivalence found in Eqs. (12) and (10). Moreover, since the matrices have binary entries we put that as constraints. This leads to the following equivalent formulation,
@@
$A_{G}PA_{H}+A_{G}Q+A_{H}R+S=2B\ .$ (14)
The optimization problem can thus be written as:
@@
We then turn the quadratic constraints (i.e. terms corresponding to $A_{G}PA_{H}$, $A_{H}R$, $PS$ and $RQ$) into linear constraints by introducing extra variables. For each term, we first replace each product $xz$ of binary variables $x$ and $z$ with a single binary variable $y$. We then add the following constraints $(1)\,y\leq z,\;(2)\,y\leq x,\;(3)\,y\geq x+z-1$. One thing to keep in mind is that $A_{G}$ is a constant, so quadratic terms with $A_{G}$ do not need to be linearized. We replace each product of the form $P[a,b]A_{H}[b,j]$ and $A_{H}[i,a]R[a,j]$ with new corresponding variables $Z^{P}_{abj}$ and $Z^{R}_{aij}$, respectively. For terms of the form $PS$ and $RQ$ we introduce new variables $Z^{PS}$ and $Z^{RQ}$, respectively. Applying these changes to Eq. (15) we obtain our ILP:
@@
Using the ILP, we can thus find LC-equivalent graphs to a given graph that have minimum number of edges. We note here that better way of solving the edge-minimization problem using ILP might exist that could allow us to go beyond graph states with 16 qubits.
Appendix C Choice of neighboring states in SA
In our implementation of SA, the state $s_{k}$ at the $k$’th iteration is given by a graph. A natural choice for the set $S(s_{k})$ of neighboring states is the set of graphs that differ from $s_{k}$ by a single local complementation. This set could have a maximum size equal to the number of vertices ($n$) in the graph. However, sometimes there are vertices where local complementation does not have any effect. Moreover, there could also be vertices that lead to isomorphic graphs after applying local complementation, as one could observe in the work of Adcock et al. [34]. If the neighboring states are chosen at random, these vertices could affect the performance of the SA algorithm. Therefore, we use a graph metric called the local clustering coefficient  [39] to guide the SA algorithm to decide where local complementation should be applied, thereby improving the efficiency of SA algorithm.
The local clustering coefficient $C_{G}(v)$ of a vertex of a graph (which will always be clear from the context) provides information on the connectedness of its neighborhood. Specifically, it is the ratio between the number of edges connecting the neighbors of the given vertex and the total number of possible edges. This metric has a maximum value of 1 (when the neighborhood is fully connected) and minimum value of zero (when the neighbors share no other edges). That is,
$C_{G}(v)=\frac{\text{Number of edges between neighbors}}{\text{Possible number connections between neighbors}}.$ (17)
@@ Appendix C Choice of neighboring states in SA
Therefore, applying a local complementation to a vertex $v$ with $C_{G}(v)\geq\frac{1}{2}$ (i.e. a more connected neighborhood) changes the local clustering coefficient to be smaller than $\frac{1}{2}$ (i.e. a less connected neighborhood). Thus, applying local complementations to vertices with a high value of $C_{G}(v)$ reduces the local edge density. Moreover, to bias vertices that have a larger number of neighbors, we multiply $C_{G}(v)$ with the degree of the vertex, leading to the heuristic quantity $M_{v}\equiv C_{G}(v)D_{v}$. Based on this heuristic, we decide on the set of states $S(s_{k})$ from which the SA algorithm could choose the potential solution $s_{\textrm{pot}}$. Thus, we bias SA into choosing vertices that have a higher value of $M_{v}$.
Figure 5: Here we plot the average edges in the initial graphs (blue line) and the edges obtained from SA using metric $M_{v}$ (orange) and random vertex choice (green). In the random vertex choice method, a potential solution $s_{\textrm{pot}}$ is found by applying local complementation to a vertex chosen uniformly at random. Initial edges are the edges in the input ER graph. For this run, $k_{\textrm{max}}$ and $T(1)$ were both set to 100.
Figure 6: (a) shows the order of fusions for constructing an RGS with 8 qubits following the Intermediate-LC protocol. This order can be used to derive the fusion network shown in (b). The fusion network enables the tracking of the LC-gates relative to fusions for each resource state. Applying LC-gates to resource states in the beginning, as shown in the fusion network, allows the fusions to be reordered flexibly. This reordering can be optimized to reduce the resource requirements. An example of such reordering is the creation of the graph $H$+$L$, as shown in (c).
@@
Appendix D Generalized repeater graph states
Here, we discuss our method for creating all-photonic generalized repeater graph states (gRGS). We consider using single photon sources and optical fusions [38] to create the desired graph state. To find the optimal number of resources required for creating gRGS, we use the protocol from Ref. [28]. We combine our algorithms with their work to find an improvement in the number of required fusions and the number of resource states (3-qubit GHZ states) for efficiently creating gRGS.
A gRGS has an $n$-vertex central graph $G$ with leaves $L$ connected to each vertex, as shown in Fig. 3(b)—we will refer to this graph as $G$+$L$. Using the SA algorithm, we can reduce the number of edges in the graph $G$ by finding an approximate MER $H$. Thus, one way to create a gRGS is the Intermediate-LC protocol where we first create the graph $H$, followed by the application of LC-gates to convert $H$ to $G$. The leaves are then fused to the central graph $G$ to create $G$+$L$.

### v2

@@ 4 Creation of generalised repeater graph states
To demonstrate the utility of our algorithms, we consider the generation of photonic repeater graph states for long distance quantum communication. Repeater graph states (RGS) [9] are a key component of all-photonic quantum repeaters, enabling the distribution of a single unit of bipartite entanglement, known as an ebit over long distances. The implementation of such all-photonic repeaters has been extensively studied [29, 30, 56, 57] but often comes with a large overhead in the number of required single photon sources. A more general class of all-photonic states called generalised repeater graph states (gRGS) was introduced in Ref. [14], which offers a better scaling between photon overhead and the number of shared ebits. Here, we apply our MER algorithms to further minimise the number of single photons and optical fusion operations for the creation of gRGS.
@@
Figure 4: (a) and (b) show the required number of resource states and fusions, respectively, for creating a single instance of GG+$L$ gRGS with 64 edges and 36 vertices (see Fig. 3(b)). We compare the naive approach of direct gRGS construction (blue curve) with the Commute-LC protocol (orange curve). The Commute-LC protocol requires nearly two orders of magnitude fewer resources than the naive approach, especially at lower fusion success probabilities. The savings decrease to less than an order of magnitude for relatively higher success probabilities.
@@
The probabilistic nature of fusions necessitates a significant amount of resources to construct practically useful graph states. Creating all-photonic graph states involves multiple rounds of fusions starting with resource states that are fused to create intermediate graph states, which are incrementally fused to construct the final graph state. For a given graph, Ref. [31] constructs a so-called fusion network, indicating the order in which fusions should occur. In general, it is desirable to fuse two graph states of similar sizes to mitigate the effects caused by fusion failure. Moreover, multiple fusion networks may exist for a given graph, each potentially requiring different amounts of resources. Ref. [31] addresses these challenges and reduces the resource overhead for photonic graph state creation by optimising (1) the number of required fusions, (2) the number of resource states, and (3) the fusion order. Moreover, they provide a numerical tool for this purpose called the OptGraphState.
@@
Here, we exploit the structure of generalised repeater graph states to optimise their experimental creation. Such graphs consist of a densely-connected central graph ($G$) and leaves ($L$) that are attached to each vertex of $G$. We find an approximate MER ($H$) of the central graph $G$ using our EDM-SA algorithm. Naively, one could fuse resource states to create GG+$L$ directly. However, a more resource-efficient method for creating a gRGS could be: (1) creating the graph $H$ using fusions, (2) applying LC-gates to convert the graph $H$ to $G$, and (3) adding leaves $L$ to $G$. We will refer to this way of construction as Intermediate-LC. However, the addition of leaves at the end constrains the order of fusions which could affect the probability of creation of the gRGS. We propose commuting the LC-gates to the beginning of the protocol that lifts this constraint and allows for an optimal fusion order. We refer to this protocol as Commute-LC. In conclusion, the Commute-LC protocol has two steps: (1) apply the LC-gates to the resource states at the beginning of the protocol, (2) follow fusion order to create the graph HH+$L$ using fusions. The LC-gates here correspond to the conversion of the graph $H$ to $G$. We should note that fusions involve the application of single-qubit Pauli correction operations, but they can be commuted through Clifford gates since these are Pauli corrections.
@@
We illustrate the construction of GG+$L$ using the Commute-LC protocol in Fig. 3(a). Following the steps in Commute-LC, we first apply LC-gates to the appropriate qubits, followed by fusing states according to the order obtained for creating HH+$L$. Applying LC-gates converts the resource states into LC-equivalent non-graph states. The Commute-LC protocol fuses these non-graph states following the order of HH+$L$ to obtain the desired graph state GG+$L$. Appendix D details how to determine locations of LC-gates and how following the order for HH+$L$ constructs the GG+$L$ graph. We note here that Ref. [31] also applies LC-gates to their qubits during their optimisation. However, these LC-gates only convert a fully-connected sub-graph (clique) into a star. In contrast, by applying EDM-SA to the central graph $G$ in Protocols 1 and 2, we generalise the use of LC-gates to simplify the fusion process not just for cliques but for arbitrary densely connected sub-graphs.
@@
In Fig. 4, we compare two methods: (1) the naive approach of creating GG+$L$ directly, and (2) creation following the Commute-LC protocol. We use the Python package OptGraphState to estimate the required resources. We consider a gRGS that can share 3 ebits across a network  [^2] Similar to Ref. [14], shown in Fig. 3(b). This gRGS has a central graph G{G} with 18 vertices and 46 edges, shown in Fig. 3(b). We find an approximate MER of this graph with 23 edges using the EDM-SA algorithm (with $k_{\textrm{max}}=10000$ and $T(1)=100$). We observe that commuting LC-gates to the beginning and following fusion order for creating HH+$L$ significantly reduces the required resources. This reduction is nearly two orders of magnitude for fusions in high-loss conditions.
5 Outlook
@@
We have proposed three algorithms to address the edge-minimisation problem. First, a simulated annealing (EDM-SA) approach efficiently finds approximate MERs for graphs with up to 100 qubits. Second, we introduce an integer linear programming (EDM-ILP) formulation that yields exact solutions, albeit with potentially exponential runtime. Third, we combine both methods into a hybrid EDM-SAILP algorithm, enabling exact MER identification for graphs with up to 16 qubits.
@@
We benchmark all algorithms on two graph models: bounded-degree and Erdős-Rényi. Additionally, we extend the ILP to handle weighted-edge minimisation and prove that this variant is NP-complete. Importantly, we show that the scalable EDM-SA approach manages to find approximate MERs with less than 10% more edges on average than the actual MERs for a large set of randomly sampled graphs. Moreover, the performance of EDM-SA could still be improved by tailoring the initial temperature and maximum iterations for specific problem instances.
@@
We have applied our EDM-SA algorithm to minimise the required resources for all-photonic generalised repeater graph states (gRGS) demonstrating its practical utility for long distance quantum communication. Specifically, we have proposed to apply LC-gates corresponding to local complementation to the resource states before fusing them to create the gRGS states. We then demonstrate that EDM-SA can be combined with existing methods from Ref. [31] to reduce the required number of single-photon sources and fusion operations by more than an order of magnitude for realistic lossy implementations.
@@
Our results open up a number of future directions. Finding the MER can reduce the number of required CZ gates for creating a desired graph state between qubits following the idea in Ref. [34]. However, there are examples of graph states where the minimum number of edges up to local complementations is greater than the minimum number of two-qubit gates needed to create the state [59, 60, 61]. It would be interesting to further investigate the exact connection between MERs and the minimal number of CZ-gates for generating qubit graph state generation
@@
Furthermore, we have focused exclusively on transformations between graph states that are allowed with local Clifford unitaries. In practical scenarios, however, one might consider arbitrary local unitaries (LU), which define a strictly coarser equivalence relation on graphs than the LC-equivalence relation, as was first shown in [62]. Graph states under local unitary equivalence are significantly less well-understood, but this topic has seen recent investigations [63, 64, 65], closely relating this equivalence to the Clifford hierarchy. One natural follow-up question is whether edge minimisation of graphs under local unitaries can similarly be phrased as an ILP (using, for example, the algorithm from [65]), and, if so, to what extent this coarser equivalence can reduce the edge count.
Acknowledgements
@@
We thank Tim Coopmans and Sebastiaan Brand for useful conversations, and Luise Prielinger for her help with the HPC. JH acknowledges funding from the Dutch Research Council (NWO) through a Veni grant (grant No.VI.Veni.222.331). JH and HS acknowledge funding from the Quantum Software Consortium (NWO Zwaartekracht Grant No.024.003.037). JB acknowledges support from The AWS Quantum Discovery Fund at the Harvard Quantum Initiative. FR acknowledges support from NSF (NSF grant No. ERC-1941583). Part of this work was performed while HS was on a research visit at Harvard University funded by a Quantum Delta NL travel grant.
@@
The package is publicly available on PyPI and can be installed with pip install graphstate-opt. Moreover, a release of our code is also available at [66] and the accompanying data at [67], with the expectation that it will serve as a valuable resource for the quantum information community in addressing practical optimisation challenges related to graph states. The usage guidelines and notes can be found at the GitHub repository https://github.com/arr0w-hs/graph_state_optimization.
@@ Appendix A LC-equivalence algorithm for graph states
$M^{T}\Omega M=\Omega~\textrm{, ~where }\Omega=\begin{bmatrix}0&I\\ I&0\\ \end{bmatrix}\$, (6)
@@
[…]
@@
Eq. (12) is a set of $n^{2}$ equations in $4n$ variables $(p_{i},q_{i},r_{i},s_{i})$ with a constraint (Eq. (10)) that is non-linear, so in principle not easily solvable. However, through an ingenious argument, Bouchet [39] proved that only a small number of solutions to Eq. 12 have to be checked to verify if there exists one that also satisfies Eq. (10), making the total runtime polynomial [39]. However for edge minimisation, we have unfortunately not been able to exploit this extra structure with our ILP (see the main text). As such, we just use Eqs. 12 and 10 to construct our ILPs.
Appendix B Derivation of EDM-ILP
In this section, we show how we formulate an ILP for edge minimisation inspired by Bouchet’s algorithm.
Let us first formally state our optimisation problem. The input is a single graph $G$ and the output is an LC-equivalent graph $H$ with a minimal number of edges. Mathematically it is framed as follows:
@@ Appendix B Derivation of EDM-ILP
We now utilise the formulation of LC-equivalence found in Eqs. (12) and (10). Moreover, since the matrices have binary entries we put that as constraints. This leads to the following equivalent formulation,
@@
$A_{G}PA_{H}+A_{G}Q+A_{H}R+S=2B\ .$
The optimisation problem can thus be written as:
@@
We then turn the quadratic constraints (i.e. terms corresponding to $A_{G}PA_{H}$, $A_{H}R$, $PS$ and $RQ$) into linear constraints by introducing extra variables. For each term, we first replace each product $xz$ of binary variables $x$ and $z$ with a single binary variable $y$. We then add the following constraints $(1)\,y\leq z,\;(2)\,y\leq x,\;(3)\,y\geq x+z-1$. One thing to keep in mind is that $A_{G}$ is a constant, so quadratic terms with $A_{G}$ do not need to be linearised. We replace each product of the form $P[a,b]A_{H}[b,j]$ and $A_{H}[i,a]R[a,j]$ with new corresponding variables $Z^{P}_{abj}$ and $Z^{R}_{aij}$, respectively. For terms of the form $PS$ and $RQ$ we introduce new variables $Z^{PS}$ and $Z^{RQ}$, respectively. Applying these changes to Eq. (14) we obtain our EDM-ILP:
@@
Using the ILP, we can thus find LC-equivalent graphs to a given graph that have minimum number of edges. We note here that better ways of solving the edge-minimisation problem using ILP might exist that could allow us to go beyond graph states with 16 qubits.
Appendix C Choice of neighbouring states in EDM-SA
In our implementation of SA, the state $s_{k}$ at the $k$’th iteration is given by a graph. A natural choice for the set $S(s_{k})$ of neighbouring states is the set of graphs that differ from $s_{k}$ by a single local complementation. This set could have a maximum size equal to the number of vertices ($n$) in the graph. However, sometimes there are vertices where local complementation does not have any effect. Moreover, there could also be vertices that lead to isomorphic graphs after applying local complementation, as one could observe in the work of Adcock et al. [37]. If the neighbouring states are chosen at random, these vertices could affect the performance of the SA algorithm. Therefore, we use a graph metric called the local clustering coefficient  [42] to guide the SA algorithm to decide where local complementation should be applied, thereby improving the efficiency of SA algorithm.
The local clustering coefficient $C_{G}(v)$ of a vertex of a graph (which will always be clear from the context) provides information on the connectedness of its neighbourhood. Specifically, it is the ratio between the number of edges connecting the neighbours of the given vertex and the total number of possible edges. This metric has a maximum value of 1 (when the neighbourhood is fully connected) and minimum value of zero (when the neighbours share no other edges). That is,
$C_{G}(v)=\frac{\text{Number of edges between neighbours}}{\text{Possible connections between neighbours}}.$ (16)
@@ Appendix C Choice of neighbouring states in EDM-SA
Therefore, applying a local complementation to a vertex $v$ with $C_{G}(v)\geq\frac{1}{2}$ (i.e. a more connected neighbourhood) changes the local clustering coefficient to be smaller than $\frac{1}{2}$ (i.e. a less connected neighbourhood). Thus, applying local complementations to vertices with a high value of $C_{G}(v)$ reduces the local edge density. Moreover, to bias vertices that have a larger number of neighbours, we multiply $C_{G}(v)$ with the degree of the vertex, leading to the heuristic quantity $M_{v}\equiv C_{G}(v)D_{v}$, similar to [69]. Based on this heuristic, we decide on the set of states $S(s_{k})$ from which the SA algorithm could choose the potential solution $s_{\textrm{pot}}$. Thus, we bias SA into choosing vertices that have a higher value of $M_{v}$.
Figure 5: Here we plot the average edges in the initial graphs (blue line) and the edges obtained from EDM-SA using metric $M_{v}$ (orange) and random vertex choice (green). In the random vertex choice method, a potential solution $s_{\textrm{pot}}$ is found by applying local complementation to a vertex chosen uniformly at random. Initial edges are the edges in the input ER graph. For this run, $k_{\textrm{max}}$ and $T(1)$ were both set to 100.
Figure 6: (a) shows the order of fusions for constructing an RGS with 8 qubits following the Intermediate-LC protocol. This order can be used to derive the fusion network shown in (b). The fusion network enables the tracking of the LC-gates relative to fusions for each resource state. To track each fusion operation, we label each fusion with a number next to it. These numbers allow us to keep track of the fusion operations after we commute the LC-operations through the fusions. Applying LC-gates to resource states in the beginning, as shown in the fusion network, allows the fusions to be reordered flexibly. This reordering can be optimised to reduce the resource requirements. An example of such reordering is the creation of the graph HH+$L$, as shown in (c).
@@
Appendix D Generalised repeater graph states
Here, we discuss our method for creating all-photonic generalised repeater graph states (gRGS). We consider using single photon sources and optical fusions [41] to create the desired graph state. To find the optimal number of resources required for creating gRGS, we use the protocol from Ref. [31]. We combine our algorithms with their work to find an improvement in the number of required fusions and the number of resource states (3-qubit GHZ states) for efficiently creating gRGS.
A gRGS has an $n$-vertex central graph $G$ with leaves $L$ connected to each vertex, as shown in Fig. 3(b)—we will refer to this graph as GG+$L$. Using the EDM-SA algorithm, we can reduce the number of edges in the graph $G$ by finding an approximate MER $H$. Thus, one way to create a gRGS is the Intermediate-LC protocol where we first create the graph $H$, followed by the application of LC-gates to convert $H$ to $G$. The leaves are then fused to the central graph $G$ to create GG+$L$.